OpenStudy (anonymous):

solve the following equation for the variable x 1/x-1 + 7/x+1 = 4x/x^2-1

OpenStudy (anonymous):

=>(x+1+x-1)/x^2-1 =4x/x^2-1 => 2x-2=4x =>2x=-2 =>x=-1

OpenStudy (anonymous):

1(x+1)+7(x-1) 4x ------------- = ---------------- x^2-1 x^2-1

OpenStudy (anonymous):

x+1+7x-7=4x 8x-4x=6 4x=6 x=3/2